OpenStudy (anonymous):

what is the sum of the series- (4/3)+(4/3^2)-(4/3^3)+(4/3^4) ...

OpenStudy (anonymous):

a=4/3 , r= 1/3 sum= (4/3) / ( 1- 1/3) = (4/3)/ (2/3) = 2

OpenStudy (anonymous):

but apparently the answers supposed to be 5/3 ...?

OpenStudy (anonymous):

ohh, but ours for some reason has subtraction as well.

OpenStudy (anonymous):

you mean the first term and third term and so on are negative???

OpenStudy (anonymous):

no, the third, fifth, etc.

OpenStudy (anonymous):

a=4/3 , r= -1/3 sum= (4/3) / ( 1+ 1/3) = (4/3)/ (4/3) = 1

OpenStudy (anonymous):

mmm...i got better solution

OpenStudy (anonymous):

4/3 + [(4/9)/ (1+4/3)]